# Multi-FAct: Assessing Multilingual LLMs’ Multi-Regional Knowledge using FActScore

### v2

@@ 3.1 Does the Factuality Differ Across Languages?
Table 1: Overall FActScore by language with mean and standard deviation (STD). The continents with the highest and second-highest values for each language are emphasized in bold and underlined, respectively.
@@
Figure 6: Continental distribution of top 20 countries that had the highest FActScore in each language. The right way to interpret this plot is, say for German, we take the twenty most accurate presidents’ biographies and then look into which continents they belong to.
3.2 Does the Factuality Geographically Differ Across Languages?
In this section, we explore whether the factuality of models varies with the geographic distribution of the language under study. Since GPT-4 performs better in low resource language as seen in figure 3, from here on, our analysis is conducted using only the outputs from GPT-4.
@@ 3.2 Does the Factuality Geographically Differ Across Languages?
Figure 7: Fine-grained geographic distribution of languages with the highest standard deviation (Chinese, Korean, Arabic) and the least standard deviation (English) is given [^5] Please note that we discarded South Africa and Uzbekistan from this analysis because they corresponded to only elements in their respective geographic categories, while other categories represent averages of two or more countries. Details of which countries belong to which category can be found in appendix A.1. The top bar represents the average number of hallucinated facts, and the bottom one denotes correct facts. The number on top of each bar represents FActScore .
To conduct an in-depth analysis, we counted which continent the top-K FActScore for each language pertains to, thereby mapping out the distribution across continents. Figure 6 presents the findings for when K equals 20, representing the top 25 percentile of the dataset.
Across analyzed languages, America and Europe are the primary focal points, underscoring their global influence. Even in languages like Korean and Chinese, distinct from these continents, the most accurate outputs primarily relate to American and European topics. For Arabic, while the highest FActScore is linked to Africa, likely reflecting its geographical roots, the second highest pertains to America. This trend highlights a Western-centric bias in the model’s factual content distribution across languages.
3.3 Geographical Biases in Factual Precision Across Subregions
In this section, we examine how languages with notable variations in factuality across continents exhibit distinct geographical biases. Further analysis involves breaking down the continents into sub-regions to compare the number of correct facts and FActScore among three languages: Chinese and Korean, which exhibit he highest standard deviation (SD), and English, which demonstrates the lowest SD among the four continents, as detailed in Table 1 of § 3.2.
As shown in Figure 5, Chinese stands out in Eastern Asia, achieving the highest number of correct facts among all regions analyzed, indicating a pronounced bias towards its main linguistic area. Korean, also categorized in Eastern Asia, shows comparable levels of factuality in Eastern Asia to that of regions associated with America and Europe, underscoring geographical biases in the model’s factual precision. English, while displaying a bias towards North America, its main region, exhibits relatively even factuality across other regions. This analysis underscores the geographical biases in factuality, with languages showing preferential accuracy towards regions where they are predominantly spoken (Figures for the other language can be found in Figure 10 in Appendix A.3.)
3.4 Is there a correlation between FActScore across languages?
Figure 8: FActScore correlation matrix computed over all the FActScore values per language.
Figure 8 shows the language-wise distribution of FActScore grouped by continents. English, Spanish, French and German have the most correlation with each other and these languages have very similar FActScore . One interesting observation is that, Swahili, even though has very similar FActScore with English, Spanish, French, and German, (table 1) does not correlate with those languages as much. This likely reflects the fact that these two languages have slightly different knowledge distributions, which we also observed in figure 6.
4 Qualitative Analysis
Model Generations with Low FActScores
Qualitative analysis on model generations, and generations with low FActScore reveals some interesting patterns. Firstly, LLMs provide much more informative outputs in English than in other language, which is also quantified in figure 4. For example, in the biography of Park Geun-hye, the former president of South Korea, GPT-4 generates the following sentence in English: “Park Geun-hye studied electrical engineering at Sogang University in Seoul and later pursued further studies at the University of Grenoble in France” while when asked in Korean, it only mentions “She studied in France (…)”. This highlights the need to develop a metric that measures beyond simple factuality of each atomic sentence, but the informativeness, or specificity of it. Some generation failures in Korean (the outliers in Figure 5) reveal that GPT-4 suggests there’s no famous person by that name and proceeds to generate a biography template. Interestingly, the English version of GPT-4 never refuses to generate a biography of the leaders in our list.
Comparison of GPT-3.5 & GPT-4
The qualitative difference between GPT-3.5 and GPT-4 is also striking. For example, while GPT-3.5 struggles to even generate fluent and coherent texts in Bengali, GPT-4 Bengali generations are mostly fluent and accurate, however the generations were short and lacking specific details. One interesting example regarding difference between GPT-3.5 and GPT-4 comes from Arabic. In GPT-3.5, when asked about Pakistani former president Mamnoon Hussain, GPT-3.5 hallucinates a biography non-existent "ambitious young man from Saudi Arabia", while GPT-4 rightfully identifies and writes the biography of former Pakistani president. GPT-4 also refused to answer more often than GPT-3.5; For example, in Chinese, it refuses to answer the biography of Sushir Koirala and claims there is no famous person of this name, however goes on to generate a fictional biography of an Indian person (with a disclaimer). GPT-4, too, does not always refrain from answering, when it doesn’t know. For example, when asked about Ollanta Humala in Chinese, it goes on to describe Ollantay Tambo, an archeological site in Peru.
5 Discussion & Limitations
In this study, we introduce a unique approach to examining the distribution of geo-culturally based knowledge using atomic factuality in long-form LLM generation. We observe that LLMs exhibit a Western-centric bias, which remains evident even in languages from non-Western regions. This discovery prompts intriguing inquiries regarding the representation of multilingual knowledge within LLMs. For instance, do LLMs learn to represent more facts about South America in Bengali by being exposed solely to content in English or Spanish? Or, to what extent is multilingual data required to acquire factual knowledge in a specific language? However, due to the proprietary nature of models like GPT-4, answering these questions proves to be challenging. Future research should delve deeper into this issue with language models that also have open datasets. Recent studies such as Dolma Soldaini et al. (2024) and OLMo Groeneveld et al. (2024) represent a step forward in this regard, although their focus is limited to monolingual English corpora and models.
One limitation of our current study pertains to small sample bias. We examine only a single sample from each country, which may not provide a comprehensive representation, especially considering that automated FActScore sometimes yields slight overestimations or underestimations. Consequently, caution is needed when interpreting individual scores. For instance, assertions such as "In Korean, the biography of the prime minister of Malaysia is 5% more accurate than that of the Romanian president" should be made with caution. In this paper, we present all findings as averages across languages, continents, or regions to mitigate sample bias. Future research should include multiple individuals, not only political leaders, from each country to alleviate this limitation.
Another limitation of our methodology is the varying durations national leaders have been in power, potentially biasing internet corpora in their favor. One solution could be to prompt models to generate a “country biography” or “capital biography” (e.g., “Write me an essay on Addis Ababa”). However, we opted not to pursue this in our study, as the original FActScore research did not include a human evaluation baseline for non-biographical domains.
6 Related Work
LLM Factuality Evaluation
The assessment of factual accuracy in natural language texts predates the advent of LLMs Guo et al. (2022). Our Multi-FAct pipeline follows the previous fact evaluation research, which involves validating claims based on external resource such as Wikipedia article Thorne et al. (2018); Krishna et al. (2022); Zhong et al. (2019) or Google search Chern et al. (2023). Our work is heavily inspired and based on Min et al. (2023), which suggests to measure the factuality of a long-form generated output of LLMs by breaking it down into atomic facts. Another line of work focuses on evaluating factual accuracy solely based on model output or internal states, eliminating the need for external knowledge bases like Wikipedia Manakul et al. (2023); Azaria and Mitchell (2023); Dhuliawala et al. (2023). While such methods offer resource efficiency, they often sacrifice accuracy compared to approaches leveraging external resources for fact verification.
Multilingual Factuality Evaluation
In the realm of multilingual factuality, Gupta and Srikumar (2021) presents the X-Fact dataset for multilingual fact-checking. Similarly, Thorne et al. (2018) have developed frameworks for multilingual fact-extraction and verification evaluation. Additionally, Aharoni et al. (2022) introduce mFACE, a dataset aimed at evaluating multilingual factual consistency. While these contributions significantly advance the understanding of multilingual factuality, they primarily focus on the assessment of factual accuracy in existing articles or benchmarks rather than examining the factual accuracy in the generative outputs of models. In this vein, Kang et al. (2024) explore the efficacy of existing metrics designed to detect hallucinations of LLM-generations, proposed for English texts, in a multilingual setting and point out they fall short in multilingual contexts. Our study extends the endeavor of developing a factuality evaluation metric that can be applied in multiple languages by introducing a Multi-Fact pipeline.
Geo-culture biases of LLMs
There has been considerable research regarding cultural and linguistic bias in LLMs Hovy and Yang (2021); Cao et al. (2023); Hershcovich et al. (2022); Huang and Yang (2023); Jin et al. (2023). A concurrent work Mirza et al. (2024) examines geographic knowledge gaps in LLMs and finds that GPT models show higher factuality in the Global North over the Global South, consistent with our findings. Manvi et al. (2024) also reach a similar conclusion. However, none of the studies concerns free-form generation with non-English languages.
7 Conclusion & Future Work
In this paper, we introduce a simple pipeline to evaluate factuality of LLMs in multilingual settings and demonstrate its effectiveness with the national leaders’ biographies in 80 countries around the world. Our findings reveal that multilingual models, such as GPT-4, tend to produce more factual content in high-resource languages, with notably longer outputs in English. Additionally, we observe a pervasive Western-centric bias across all languages; on average, even Asian and African languages demonstrate better performance for American and European leaders. However, a more detailed geographical analysis indicates that languages still reflect some aspects of their local geographic distribution, albeit to a lesser degree.
While factuality might be an interesting way to look into a model’s knowledge about a certain person, our current implementation treats every fact equally. FActScore does not differentiate between the informational value of facts. For instance, the statements “Person X attended MIT” and “Person X attended a famous university” are treated as equivalent, though the former is obviously more informative and useful. Future research should aim to distinguish between specific, valuable facts and generic, less informative ones.
Another area for future investigation could involve examining the consistency of model-generated facts across different languages; for example, whether the model’s depiction of Barack Obama’s childhood in English aligns with its portrayal in Chinese, or whether it keeps its focus solely on his political endeavors.
Appendix A Appendix
A.1 List of Countries and Presidents
@@
Table 2 lists all the names of the countries, presidents and their Geographic locations. Please note that, we changed Australia and New Zealand to Oceania in the UN geoscheme for plotting purpose.
@@
Table 2:
A.2 Multi-FAct Implementation Details
We replicated the original FActScore implementation and replaced InstructGPT for atomic fact generation with Mistral-7B. We also used Mistral 7B for the retrieval-based verification step. The final version uses retrieval + Mistral-7B + NPM. NPM threshold was set to 0.3.
A.2.1 Replication of original FActScore
We provide a comparison between our implementation with the results reported in the paper. For all the following experiments, we used a subset of ChatGPT-generation human-annotated factual data from the original paper.
Metric Human FS FS (ours) Error (ours) Error (Original)
Value 0.626 $\pm$ 0.238 0.637 $\pm$ 0.207 -0.011 $\pm$ 0.129 -0.0383 $\pm$ 0.121
Table 3: Summary of Replication. Error refers to the difference between human FActScore and FActScore determined by our method. Error (Original) refers to the error by the best-performing method reported on the paper. We also report the standard deviation alongside each metric.
Note that the Human FActScore value in the table 3 is different from what is provided in the paper as we averaged over our subset only. We replicated the best-performing method (retrieval + NPM) on this dataset reported in the paper and marked it as Error (Original). Our higher accuracy is likely due to Mistral 7B’s better capabilities and our better hyperparameter tuning.
Please note that there is some stochasticity in the automated evaluation of FActScore , which is why we need to be cautious of individual scores, and only work with scores averaged over a large number of samples.
A.2.2 Evaluation of Multi-FAct Pipeline
The basis of our pipeline relies crucially on the GPT3.5-based translation (Step 3 in figure 2). It is imperative to see whether this translation step hurts the automated estimation of FActScore in a multilingual setting.
Evaluation Method:
The basis of our evaluation method is that the factuality of a text is translation invariant, ie, factual knowledge in a text should not change when it is translated to other languages. To evaluate this, we first take the human-annotated examples provided in original FActScore , and translate them into four languages using GPT4. We use GPT4 as previous studies Jiao et al. (2023) revealed GPT4 to be a good translator and better than GPT3.5 and commercially available machine translators. Manual inspection by native speakers of a subset of our languages also confirmed this.
Language/Dataset FActScore
Human eval (EN) 0.626 $\pm$ 0.238
EN 0.637 $\pm$ 0.207
KOR 0.625 $\pm$ 0.211
BN 0.600 $\pm$ 0.216
FR 0.638 $\pm$ 0.200
SW 0.654 $\pm$ 0.196
Table 4: FActScore evaluation Across Languages. EN refers to FActScore evaluation on original English human annotation; other languages refer to FActScore evaluation on GPT4 translation of the human annotation.
Figure 9: Correlation of FActScore evaluated on translated human-annotated data.
We then applied our multilingual FActScore pipeline (Figure 2) on this GPT4 translated text. The premise is that if our pipeline (which involves GPT3.5 translation [^6] We did not use GPT4 translation in Multi-FAct pipeline to reduce API costs) provides similar FActScore as the original English annotated data, it might provide good evidence that having a translation step in our pipeline doesn’t hurt FActScore estimation significantly.
The results are provided in Table 4. We see that running Multi-FAct on non-English data does not significantly degrade FActScore estimation.
We also note on the bottom row of Figure (9) that the correlation between FActScore evaluated in English and FActScore evaluated in other languages for the texts are very high; and all of them are very closely correlated with human-judgement.
A.3 Additional Results
Regional Breakdown (GPT4):
Figure 10 shows the regional breakdown for the other languages that were not included in Figure 5.
(a) Spanish
(b) German
(c) Swahili
(d) French
(e) Bengali
Figure 10: Fine-grained Geographic Distribution for Remaining Languages

### v3

@@ 5.2 Does the Factuality Geographically Differ Across Languages?
Table 3: Overall FActScore breakdown of GPT4 by language with mean and standard deviation (STD). The continents with the highest and second-highest values for each language are emphasized in bold and underlined, respectively. Note that for all language except for Swahili, GPT4 is leaning towards American and European leaders in terms of factual accuracy.
6 Conclusion
In this paper, we introduce Multi-FAct , an extension of FActScore in multilingual setting to address the critical gap in factuality evaluation for non-English free-form generation from LLMs. We empirically demonstrate the viability of conducting factuality evaluation in the multilingual biography generation task. Our approach of translating generated content and comparing it against English Wikipedia proved effective, offering a scalable method for assessing factual precision across multiple languages. Additionally, we find that sufficiently comprehensive non-English sources can serve as alternatives to English sources using the same pipeline.
Our analysis of LLMs’ outputs across different languages and geographically contextualized questions reveals a consistent advantage for English in terms of both factual accuracy and the quantity of generated facts, as well as better performance for content related to North America and Europe across languages. Multi-FAct serves as a valuable tool for investigating these biases, emphasizing the need for efforts to improve the cultural and geographic fairness of factual generations from LLMs.
7 Limitations
Our work shares some of the same limitations that original FActScore faces. Throughout the whole paper, we assume English Wikipedia as the most authoritative knowledge base for fact verification. However, it is known that Wikipedia itself has a Western-centric bias Naous et al. (2023). Moreover, when combining multiple non-English Wikipedia, we do not take into account when Wikipedia articles in different languages contradict each other, which can happen for controversial topics. It is also worth noting that FActScore and Multi-FAct both do not differentiate between the informational value of facts. For example, statements like “Person X attended MIT” and “Person X attended a renowned university” are treated equally, despite the former being more specific and informative. Future research should aim to distinguish between specific, valuable facts and generic, less informative ones.
Acknowledgments
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No. RS-2024-00509258, AI Guardians: Development of Robust, Controllable, and Unbiased Trustworthy AI Technology). We thank Pang Wei Koh and Sewon Min for providing valuable feedback on an earlier draft of this work.
Appendix A Comparison with Baseline
We provide a comparison of different components of Multi-FAct below.
Trivial baselines: Random prediction: 8.86% Always "Supported": - 41.15% Always "Not Supported": 58.86%
Method Arabic Swahili Bengali Chinese Korean French German Spanish Mean
No Retrieval ChatGPT -0.268 -0.282 -0.260 -0.265 -0.277 -0.278 -0.288 -0.290 -0.276
Retrieval + ChatGPT -0.163 -0.190 -0.165 -0.168 -0.177 -0.177 -0.175 -0.187 -0.175
Retrieval + Mistral -0.150 -0.180 -0.169 -0.162 -0.172 -0.171 -0.175 -0.174 -0.169
NPM -0.097 -0.133 -0.105 -0.096 -0.098 -0.140 -0.149 -0.153 -0.121
Retrieval + Mistral + NPM 0.028 -0.016 0.022 0.029 0.006 -0.013 -0.013 -0.012 0.004
Table 4: Multi-FAct pipeline component comparison across different languages.
In Table 4:
  • ChatGPT refers to GPT3.5-turbo.
  • No Retrieval ChatGPT means simply asking ChatGPT whether a fact is correct or not (closed book setting).
  • Retrieval + ChatGPT implies retrieving relevant passages from Wikipedia and then answering whether a fact is correct or not using GPT3.5.
  • Retrieval + Mistral refers to retrieving passages and then using Mistral 7B to answer whether a fact is correct or not. This is similar to the previous setting except for the fact verification LLM is being switched to an open-source model.
  • NPM refers to Non-Parametric Model (Min et al. (2022)), which requires having access to the Wikipedia article (ie, retrieval enabled).
  • The final method is the ensemble version of NPM and Retrieval + Mistral and it performs the best (also reported in original FActScore work (Min et al. (2023))). This is the one we used throughout our paper for other analyses.
While GPT3.5 and Mistral 7B are very similar at atomic fact breakdown, we use Mistral 7B in all experiments to save cost.
Appendix B Effect of non-LLM Based Translation
In section 3.4.2, we showed that GPT3.5-based translation does not hurt the FActScore estimate. We have also explained the reason we use GPT3.5 over commercial translation systems like Google Translate is because Google Translate is not able to keep the gender of the subject consistent across the generation. Table 5 provides a comparison of translation of FActScore estimation when the translation is done by Google Translate vs GPT3.5.
Source Language GPT3.5 Error GTranslate Error
Arabic 0.028 0.051
Swahili -0.0157 0.0168
Bengali 0.0219 0.057
Chinese 0.0289 0.0883
Korean 0.0063 0.0453
French -0.0126 -0.0106
German -0.0127 -0.0019
Spanish -0.0115 0.0581
Mean 0.0041 0.0380
Table 5: FActScore estimation error rate comparison when translation to English was done using GPT-3.5 and GTranslate across different languages
The experiment set up for Table 5 was the same as Table 2. The small difference in Error with GPT3.5 column arises from the run-to-run stochasticity of the FActScore estimation.
Can we skip the translation step altogether?
We also experimented with generating atomic facts in the target language (thus, skipping the translation step) with French and Korean using GPT3.5 Turbo. Qualitative analysis with native Korean speakers revealed that the generated atomic facts are quite reasonably good. However, because there is no NPM model available for Korean and French, the FActScore estimates using atomic facts generated with those methods do not match the best-performing method (Retrieval+Mistral+NPM) provided in the above table. Moreover, tokenizer fertility for non-English languages is often much worse than for English, which makes it much more computationally expensive to do atomic fact breakdown in Korean. At the time of writing this paper, no small open-source model, like Mistral 7B, could perform atomic fact breakdown in a diverse multilingual setting.
Appendix C Additional Results
C.1 Results from other models
Table 6 shows the benchmark results from Claude-3 and Gemini-1.0-pro model. We chose to use GPT4 for most of our analysis because GPT4 has the best performance in low resource languages.
Language (Claude) haiku opus sonnet gemini-1.0-pro gpt3.5 gpt4-turbo gpt4o-mini gpt4o
English 0.68 0.70 0.62 0.44 0.57 0.62 0.54 0.58
French 0.65 0.64 0.60 0.42 0.59 0.61 0.56 0.58
Spanish 0.66 0.70 0.60 0.47 0.57 0.63 0.59 0.59
German 0.63 0.66 0.58 0.43 0.57 0.63 0.55 0.60
Chinese 0.27 0.43 0.19 0.40 0.40 0.49 0.49 0.53
Korean 0.39 0.45 0.27 0.40 0.18 0.48 0.49 0.52
Arabic 0.33 0.45 0.26 0.32 0.24 0.53 0.46 0.51
Bengali 0.26 0.36 0.18 0.40 0.12 0.58 0.48 0.58
Swahili 0.57 0.62 0.54 0.46 0.59 0.62 0.65 0.65
Mean 0.49 0.56 0.43 0.42 0.43 0.58 0.53 0.57
Table 6: FActScore estimation of the other models on the same national leaders’ biography generation task. Note that only GPT4-turbo has the overall highest performance and on average GPT4 series models have good performance across all languages.
Note that model response size varies significantly across languages and models, which is not captured in this table. Similar to what is shown in figure 6, most of the models in this table tend to generate much longer responses for English and high-resource languages. The average response length also differs from model family to model family.
C.2 Effect of Reference Length
Sampling Method
We randomly collected 10,000 people from 9 different Wikipedia using Wikipedia API and sorted them by their byte length. After that, we kept the top 100 entities that also had an English Wikipedia. Note that since we did this for each language separately, the list of 100 people differs for each language. However, manual inspection revealed that these entities are usually famous politicians, religious figures, famous athletes, and singers.
Table 7 shows the word count distribution for the selected Wikipedia articles.
Language Mean Text Length (Language Wiki) Mean Text Length (English Wiki)
Spanish 3353.94 6895.68
French 3617.13 6569.71
German 3519.03 5723.78
Korean 1520.79 5144.43
Chinese 2879.44 4937.70
Arabic 2829.48 5648.35
Swahili 865.85 5003.25
Bengali 2772.49 5829.74
Table 7: Comparison of mean text lengths (in words) in Various Language Wikipedia after translating to English vs length of the same articles English Wikipedia. For the same subjects, English Wikipedia has much larger articles than the other languages.
Appendix D Geographical Biases in Factual Precision Across Subregions
@@
Figure 7: Continental distribution of top 20 countries that had the highest FActScore in each language for GPT4. The way to interpret this plot is, say for German, we sort the twenty most accurate presidents’ biographies and then look up which continents they belong to. Note that for almost all languages, the top twenty biographies belong to subjects from America and Europe.
We examine how languages with notable variations in factuality across continents exhibit distinct geographical biases. Further analysis involves breaking down the continents into sub-regions to compare the number of correct facts and FActScore among three languages: Chinese and Korean, which exhibit he highest standard deviation (SD), and English, which demonstrates the lowest SD among the four continents, as detailed in Table 3 of § 5.2.
Chinese stands out in Eastern Asia, achieving the highest number of correct facts among all regions analyzed, indicating a pronounced bias towards its main linguistic area. Korean, also categorized in Eastern Asia, shows comparable levels of factuality in Eastern Asia to that of regions associated with America and Europe, underscoring geographical biases in the model’s factual precision. English, while displaying a bias towards North America, its main region, exhibits relatively even factuality across other regions.
Figure 8 shows the fine-grained geographic distribution of languages with the highest standard deviation (Chinese, Korean, Arabic) and the least standard deviation (English).
@@
Figure 8: Fine-grained geographic distribution of languages with the highest standard deviation (Chinese, Korean, Arabic) and the least standard deviation (English) is given. The top bar represents the average number of hallucinated facts, and the bottom one denotes correct facts. The number on top of each bar represents FActScore .
Appendix E List of Countries and Presidents
@@
Table 8 lists all the names of the countries, presidents and their Geographic locations. Please note that, we changed Australia and New Zealand to Oceania in the UN geoscheme for plotting purposes.
@@
Table 8: List of Countries and Presidents Used in Our Experiment